The position of a particle along a straight line is given by s= (1.5t^3 − 13.5t^2+ 22.5t)ft, where t is in seconds. Determine the position of the particle when t = 6 s and the total distance it travels during the 6s time interval.
1
Expert's answer
2020-08-27T10:41:04-0400

The position of a particle

"s(t)=1.5t^3-13.5t^2+22.5"

Hence

"s(6)=1.5\times 6^3-13.5\times 6^2+22.5=-139.5\:\rm ft"

The total distance that particle travels during the 6s time interval

"d=|s(6)-s(0)|=|-139.5-22.5|=-162\:\rm ft"